OpenStudy (anonymous):

Find the third term in the binomial series from (a+b)^11

OpenStudy (anonymous):

55(a^9)(b^2)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[\dbinom{11}{2}a^9b^2\]

OpenStudy (anonymous):

\[\dbinom{11}{2}=\frac{11\times 10}{2}=11\times 5=55\]

OpenStudy (across):

Recall that this follows from\[(a+b)^n=\sum_{k=0}^{n}\binom{n}{k}a^{n-k}b^{k}.\]